# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 7. Curves with a largest possible hull
The following result gives a condition for the hull to be semi-simple.
@@
(7.1) $\lim_{t\to\infty}\varphi(t).(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and all linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$. Then, the observable hull $H_{\varphi}$ of $\varphi$ is a connected semi-simple with no compact factors, and the action of $H_{\varphi}$ on $\mathbb{R}^{n}$ is irreducible.
@@
Since $\text{zcl}(H_{\varphi})$ is observable, we have a finite dimensional rational representation $V$ of $\text{SL}(n,\mathbb{R})$ and a $v\in V$ such that $\text{zcl}(H_{\varphi})$ is the isotropy group of $v$. Note that $\varphi(t).v,t\geq 0$ is bounded since $\varphi$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$. Since we assume (7.1), it follows by Lemma 2.1 that the orbit $\text{SL}(n,\mathbb{R}).v$ is Zariski closed. Then, Matsushima criterion [Matsushima_crit] implies that $\text{zcl}(H_{\varphi})$ is reductive. Since $H_{\varphi}$ is connected and generated by one-parameter unipotent subgroups, it follows that $H_{\varphi}$ is a connected semi-simple group with no compact factors. Finally, we prove that $H_{\varphi}$ acts on $\mathbb{R}^{n}$ irreducibly. Suppose for contradiction that $\mathbb{R}^{n}=V_{1}\oplus V_{2}$ where both $V_{1}$ and $V_{2}$ are non-trivial sub-spaces invariant under $H_{\varphi}$. Let $v_{1},...,v_{k}$ by a basis for $V_{1}$. Here $1\leq k<n$ because $V_{2}$ is not trivial. Since semi-simple groups have no non-trivial characters, we obtain that $H_{\varphi}.(v_{1}\wedge\cdots\wedge v_{k})=v_{1}\wedge\cdots\wedge v_{k}$. Since $\varphi([0,\infty))$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$, we get $\varphi(t).(v_{1}\wedge\cdots\wedge v_{k})$ is bounded as $t\to\infty$. This contradicts (7.1). ∎
Let $\varphi$ be the curve given by (1.5). First, we will verify that $\varphi$ satisfies the following stronger divergence property.
@@
For any $1\leq k\leq n$, and $0\neq v\in\wedge^{k}\mathbb{R}^{n+1}$, $\lim_{t\to\infty}\varphi(t).v\to\infty$ as $t\to\infty$.
@@
We denote by $e_{0},e_{1},...,e_{n}$ the canonical basis vectors of $\mathbb{R}^{n+1}$, where $e_{i}$ denotes the vectors whose coordinates equal 00 besides the $(i+1)$th position in which $1$ appears. Fix an integer $1\leq k\leq n$ and let $\mathcal{I}$ be the collection of all subsets of $\{0,\ldots,n\}$ of cardinality $k$. Let $\varphi$ be the curve given by (1.5). For any $I:=\{i_{1},\ldots,i_{k}\}\in\mathcal{I}$, where $i_{1}<i_{2}<\cdots<i_{k}$, let $e_{I}=e_{i_{1}}\wedge\cdots\wedge e_{i_{k}}$. There are two cases for $\varphi(t).e_{I}$:
@@
$\varphi(t).e_{I}=$ $f_{0}(t)e_{0}\wedge[f_{i_{2}}(t)e_{0}+h_{i_{2}}(t)^{-1}e_{i_{2}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
@@
$\varphi(t).e_{I}=$ $[f_{i_{1}}(t)e_{0}+h_{i_{1}}(t)^{-1}e_{i_{1}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
@@
Let $\mathbf{v}\in\wedge^{k}\mathbb{R}^{n+1}\setminus\{0\}$. Then $\mathbf{v}=\sum_{I\in\mathcal{I}}c_{I}e_{I}$, where $c_{I}\in\mathbb{R}$. Then,
(7.4) $\varphi(t).\mathbf{v}=\sum_{I\in\mathcal{I}}C_{I}(t)e_{I},$
@@
Let $\mathcal{I}_{\mathbf{v}}=\{I\in\mathcal{I}:c_{I}\neq 0\}\neq\varnothing$, as $\mathbf{v}\neq 0$. Let $i_{1}=\min\cup\mathcal{I}_{\mathbf{v}}$. We pick $I\subset\mathcal{I}_{\mathbf{v}}$ such that $i_{i}\in I$. Let $J=I\setminus\{i_{1}\}$. Then, $c_{J\cup\{l\}}=0$ for every $l<i_{1}$. Therefore, by (7.5)
@@
(7.11) $\rho(s):=\begin{bmatrix}1&sv\\ 0&1\end{bmatrix}\in\text{SL}(n+1,\mathbb{R}),\,\forall s\in\mathbb{R}.$
@@
Let $\{e_{0},\ldots,e_{n}\}$ denote the standard basis of $\mathbb{R}^{n+1}$. Let $(v_{1},\ldots,v_{n})\in\mathbb{R}^{n}$ be as in Lemma lem:ps group of our main example. Let $W_{1}$ denote the ortho-complement of $v=\sum_{i=1}^{n}v_{i}e_{i}$ in the span of $\{e_{1},\ldots,e_{n}\}$. Then the fixed points space of $\rho(s)$ in $\mathbb{R}^{n+1}$ is $W:=\mathbb{R}e_{0}+W_{1}$.
Now let $\beta$ denote the correcting definable curve such that $\beta(t)\varphi(t)\subset H_{\varphi}$. Since $\beta(t)\to\beta(\infty)\in\text{SL}(n+1,\mathbb{R})$ as $t\to\infty$, by (7.12), we conclude that $\rho_{1}(s):=\beta(\infty)\rho(s)\beta(\infty)^{-1}\in H_{\varphi}$ for all $s\in\mathbb{R}$. By Lemma 7.2 and Lemma 7.1, $H_{\varphi}$ is a semisimple group acting irreducibly on $\mathbb{R}^{n+1}$. Therefore, by Jacobson-Morosov theorem, there exists a homomorphism $\eta:\text{SL}(2,\mathbb{R})\to H_{\varphi}$ such that $\eta(u(s))=\rho_{1}(s)$, where $u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}\in\text{SL}(2,\mathbb{R})$. As noted earlier, the fixed points space of $\rho_{1}$ on $\mathbb{R}^{n+1}$ is $\beta(\infty)W$, and it has dimension $n$. Expressing the representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{n+1}$ via $\eta$ as a direct sum of irreducible representations of $\text{SL}_{2}(\mathbb{R})$, and noting that the dimension of the $u(s)$-fixed subspace is $n$, we conclude that $\mathbb{R}^{n+1}=V_{1}\oplus V_{2}$, where $V_{1}$ is isomorphic to the standard representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{2}$, and $\text{SL}(2,\mathbb{R})$ acts trivially on $V_{2}$. Therefore, by [shah2023equidistribution, Theorem A.7], either $H_{\varphi}=\text{SL}(n+1,\mathbb{R})$ or $n+1=2d$ for some $d\geq 2$ and $H_{\varphi}$ is conjugate to the standard symplectic group $\text{SP}(\mathbb{R}^{2d})$.
@@ Appendix A Kempf’s Lemma for real algebraic groups
For finite-dimensional real vector spaces $V$ and $W$, we say that $P:V\to W$ is a polynomial map if, after choosing bases for $V$ and $W$, the map $P$ in those coordinates is a multi-variable polynomial in each coordinate. We say that the map is $G$-equvairant if $g.P(x)=P(g.x)$ for all $g\in G$ and $x\in V$.
@@
Let $G$ be the set of $\mathbb{R}$-points of an algebraic group defined over $\mathbb{R}$. Let $V$ be a rational representation of $G$. Suppose that for $v\in V$, it holds that $S:=\text{zcl}(G.v)\smallsetminus G.v$ is non-empty. Then there exists a rational representation $W$ and an equivariant polynomial map $P:V\to W$ such that $P(S)=\{0\}$ and $P(v)\neq 0$.
@@
First, note that since $\text{zcl}(G.v)$ and $G.v$ are $G$-invariant, we get that $S$ is $G$-invariant. Let $\mathbb{R}[V]$ be the coordinate ring on the vector space $V$ and consider $I\subseteq\mathbb{R}[V]$ the polynomial ideal of the polynomials vanishing on $S$. For $d\in\mathbb{N}$, let $I(d)\subseteq I$ be the subset of polynomials in $I$ with degree bounded by $d$. We may choose $d$ such that the ideal generated by $I(d)$ is $I$. Notice that $I(d)$ is a finite dimensional vector space. Consider the right action of $G$ on $\mathbb{R}[V]$ defined by
(A.1) $(f.g)(x):=f(g.x),~{}g\in G,~{}x\in V,~{}f\in\mathbb{R}[V].$
The action preserves the degree of the polynomial. So, it preserves $I(d)$ because $S$ is $G$-invariant. Let $I(d)^{*}$ be the space of linear functionals on $I(d)$. Using the representation (A.1), we get a left action on $I(d)^{*}$ defined by
(A.2) $g.l(f):=l(f.g),~{}g\in G,~{}l\in I(d)^{*},~{}f\in I(d).$
@@
where $\text{ev}_{x}$ is the functional mapping $f\in I(d)$ to $f(x)$. It is straightforward to verify that $P$ is an equivariant polynomial map. Finally, we claim that $P(v)=\text{ev}_{v}$ is a non-zero functional. By the definition of $S$ and $P$, $P(x)=0$ for all $x\in S$. It remains to show that $P(v)\neq 0$. For that, recall the closed orbit lemma [Bo91, Section 1.8], which states that $S$ is Zariski closed. This implies that there exists a $p\in I(d)$ such that $p(S)=0$ and $p(v)\neq 0$. Since $P(v)(p)=p(v)$, $P(v)\neq 0$. ∎

### v2

@@ 7. Curves with a largest possible hull
The main goal of this section is to prove Proposition 1.9 on the equidistribution of the curve (1.5). As an overview, the key idea is that the curve (1.5) will be strongly non-contracting, in the sense that it takes nonzero wedges to $\infty$, Lemma 7.2. The following result, of independent interest, implies the hull of (1.5) is semi-simple and acts irreducibly on $\mathbb{R}^{n}$. Finally, a key classification of [shah2023equidistribution, Theorem A.7] will allow to prove that the hull equals to the full group $\text{SL}(n,\mathbb{R})$.
@@
(7.1) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and all linearly independent vectors $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$. Then, the hull $H_{\varphi}$ of $\varphi$ is a connected semi-simple with no compact factors, and the action of $H_{\varphi}$ on $\mathbb{R}^{n}$ is irreducible.
@@
Since $\text{zcl}(H_{\varphi})$ is observable, we have a finite-dimensional rational representation $V$ of $\text{SL}(n,\mathbb{R})$ and $v\in V$ such that $\text{zcl}(H_{\varphi})$ is the isotropy group of $v$. Note that $\varphi(t)\cdot v,t\geq 0$ is bounded since $\varphi$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$. Since we assume (7.1), it follows by Lemma 2.1 that the orbit $\text{SL}(n,\mathbb{R})\cdot v$ is Zariski closed. Then, Matsushima criterion [Matsushima_crit] implies that $\text{zcl}(H_{\varphi})$ is reductive. Since $H_{\varphi}$ is connected and generated by one-parameter unipotent subgroups, it follows that $H_{\varphi}$ is a connected semi-simple group with no compact factors. Finally, we prove that $H_{\varphi}$ acts on $\mathbb{R}^{n}$ irreducibly. Suppose for contradiction that $\mathbb{R}^{n}=V_{1}\oplus V_{2}$ where both $V_{1}$ and $V_{2}$ are nontrivial sub-spaces invariant under $H_{\varphi}$. Let $v_{1},\dots,v_{k}$ be a basis for $V_{1}$. Here $1\leq k<n$ because $V_{2}$ is not trivial. Since semi-simple groups have no nontrivial characters, we obtain that $H_{\varphi}\cdot(v_{1}\wedge\cdots\wedge v_{k})=v_{1}\wedge\cdots\wedge v_{k}$. Since $\varphi([0,\infty))$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$, we get $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})$ is bounded as $t\to\infty$. This contradicts (7.1). ∎
Let $\varphi$ be the curve given by (1.5). We now verify that $\varphi$ satisfies a divergence property, which is stronger than (7.1).
@@
For any $1\leq k\leq n$, and $0\neq v\in\bigwedge^{k}\mathbb{R}^{n+1}$, $\lim_{t\to\infty}\varphi(t)\cdot v\to\infty$ as $t\to\infty$.
@@
We denote by $e_{0},e_{1},\dots,e_{n}$ the canonical basis vectors of $\mathbb{R}^{n+1}$, where $e_{i}$ denotes the vectors whose coordinates equal 00 besides the $(i+1)$th position in which $1$ appears. Fix an integer $1\leq k\leq n$ and let $\mathcal{I}$ be the collection of all subsets of $\{0,\ldots,n\}$ of cardinality $k$. Let $\varphi$ be the curve given by (1.5). For any $I:=\{i_{1},\ldots,i_{k}\}\in\mathcal{I}$, where $i_{1}<i_{2}<\cdots<i_{k}$, let $e_{I}=e_{i_{1}}\wedge\cdots\wedge e_{i_{k}}$. There are two cases for $\varphi(t)\cdot e_{I}$:
@@
$\varphi(t)\cdot e_{I}=$ $f_{0}(t)e_{0}\wedge[f_{i_{2}}(t)e_{0}+h_{i_{2}}(t)^{-1}e_{i_{2}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
@@
$\varphi(t)\cdot e_{I}=$ $[f_{i_{1}}(t)e_{0}+h_{i_{1}}(t)^{-1}e_{i_{1}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
@@
Let $\mathbf{v}\in\bigwedge^{k}\mathbb{R}^{n+1}\setminus\{0\}$. Then $\mathbf{v}=\sum_{I\in\mathcal{I}}c_{I}e_{I}$, where $c_{I}\in\mathbb{R}$. Then,
(7.4) $\varphi(t)\cdot\mathbf{v}=\sum_{I\in\mathcal{I}}C_{I}(t)e_{I},$
@@
Let $\mathcal{I}_{\mathbf{v}}=\{I\in\mathcal{I}:c_{I}\neq 0\}\neq\varnothing$, as $\mathbf{v}\neq 0$. Let $i_{1}=\min\cup\mathcal{I}_{\mathbf{v}}$. We pick $I\subseteq\mathcal{I}_{\mathbf{v}}$ such that $i_{i}\in I$. Let $J=I\setminus\{i_{1}\}$. Then, $c_{J\cup\{l\}}=0$ for every $l<i_{1}$. Therefore, by (7.5)
@@
(7.11) $\rho(s):=\begin{bmatrix}1&sv\\ 0&I_{n}\end{bmatrix}\in\text{SL}(n+1,\mathbb{R}),\,\forall s\in\mathbb{R}.$
@@
Let $\{e_{0},\ldots,e_{n}\}$ denote the standard basis of $\mathbb{R}^{n+1}$. Let $(v_{1},\ldots,v_{n})\in\mathbb{R}^{n}$ be as in Lemma 7.3. Let $W_{1}$ denote the orthogonal complement of $v=\sum_{i=1}^{n}v_{i}e_{i}$ in the span of $\{e_{1},\ldots,e_{n}\}$. Then the fixed points space of $\rho(s)$ in $\mathbb{R}^{n+1}$ is $W:=\mathbb{R}e_{0}+W_{1}$.
Now let $\beta$ denote the correcting definable curve such that $\beta(t)\varphi(t)\subseteq H_{\varphi}$. Since $\beta(t)\to\beta(\infty)\in\text{SL}(n+1,\mathbb{R})$ as $t\to\infty$, by (7.12), we conclude that $\rho_{1}(s):=\beta(\infty)\rho(s)\beta(\infty)^{-1}\in H_{\varphi}$ for all $s\in\mathbb{R}$. By Lemma 7.2 and Lemma 7.1, $H_{\varphi}$ is a semisimple group acting irreducibly on $\mathbb{R}^{n+1}$. Therefore, by Jacobson-Morosov theorem, there exists a homomorphism $\eta:\text{SL}(2,\mathbb{R})\to H_{\varphi}$ such that $\eta(u(s))=\rho_{1}(s)$, where $u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}\in\text{SL}(2,\mathbb{R})$. As noted earlier, the fixed points space of $\rho_{1}$ on $\mathbb{R}^{n+1}$ is $\beta(\infty)W$, and it has dimension $n$. Expressing the representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{n+1}$ via $\eta$ as a direct sum of irreducible representations of $\text{SL}_{2}(\mathbb{R})$, and noting that the dimension of the $u(s)$-fixed subspace is $n$, we conclude that $\mathbb{R}^{n+1}=V_{1}\oplus V_{2}$, where $V_{1}$ is isomorphic to the standard representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{2}$, and $\text{SL}(2,\mathbb{R})$ acts trivially on $V_{2}$. Therefore, by [shah2023equidistribution, Theorem A.7], either $H_{\varphi}=\text{SL}(n+1,\mathbb{R})$ or $n+1=2d$ for some $d\geq 2$ and $H_{\varphi}$ is conjugate to the standard symplectic group $\text{SP}(\mathbb{R}^{2d})$.
@@ Appendix A Kempf’s Lemma for real algebraic groups
For finite-dimensional real vector spaces $V$ and $W$, we say that $P:V\to W$ is a polynomial map if, after choosing bases for $V$ and $W$, the map $P$ in those coordinates is a multi-variable polynomial in each coordinate. We say that the map is $G$-equvairant if $g\cdot P(x)=P(g\cdot x)$ for all $g\in G$ and $x\in V$.
@@
Let $G$ be the set of $\mathbb{R}$-points of an algebraic group defined over $\mathbb{R}$. Let $V$ be a rational representation of $G$. Suppose that for $v\in V$, it holds that $S:=\text{zcl}(G\cdot v)\setminus G\cdot v$ is non-empty. Then there exists a rational representation $W$ and an equivariant polynomial map $P:V\to W$ such that $P(S)=\{0\}$ and $P(v)\neq 0$.
@@
First, note that since $\text{zcl}(G\cdot v)$ and $G\cdot v$ are $G$-invariant, we get that $S$ is $G$-invariant. Let $\mathbb{R}[V]$ be the coordinate ring on the vector space $V$ and consider $I\subseteq\mathbb{R}[V]$ the polynomial ideal of the polynomials vanishing on $S$. For $d\in\mathbb{N}$, let $I(d)\subseteq I$ be the subset of polynomials in $I$ with degree bounded by $d$. We may choose $d$ such that the ideal generated by $I(d)$ is $I$. Notice that $I(d)$ is a finite-dimensional vector space. Consider the right action of $G$ on $\mathbb{R}[V]$ defined by
(A.1) $(f.g)(x):=f(g\cdot x),~{}g\in G,~{}x\in V,~{}f\in\mathbb{R}[V].$
The action preserves the degree of a polynomial. So, it preserves $I(d)$ because $S$ is $G$-invariant. Let $I(d)^{*}$ be the space of linear functionals on $I(d)$. Using the representation (A.1), we get a left action on $I(d)^{*}$ defined by
(A.2) $g\cdot l(f):=l(f\cdot g),~{}g\in G,~{}l\in I(d)^{*},~{}f\in I(d).$
@@
where $\text{ev}_{x}$ is the functional mapping $f\in I(d)$ to $f(x)$. It is straightforward to verify that $P$ is an equivariant polynomial map. Finally, we claim that $P(v)=\text{ev}_{v}$ is a nonzero functional. By the definition of $S$ and $P$, $P(x)=0$ for all $x\in S$. It remains to show that $P(v)\neq 0$. For that, recall the closed orbit lemma [Bo91, Section 1.8], which states that $S$ is Zariski closed. This implies that there exists a $p\in I(d)$ such that $p(S)=0$ and $p(v)\neq 0$. Since $P(v)(p)=p(v)$, $P(v)\neq 0$. ∎
Appendix B Basic notions in the theory of o-minimal structures
In order to keep our paper self contained, we will now recall the basic definitions and results on o-minimal structures on the field of real numbers. We refer to [van_den_dries_omin_book] and [Geom_cat_and_o_min_str] for more details.
We begin with the notion of a structure on the ordered field of real numbers, and later introduce the stronger notion of an o-minimal structure.
Definition B.1.
A structure $\mathscr{S}$ on the ordered field of real numbers is a sequence of families of sets
$\mathscr{S}_{d}\subseteq\{\text{subsets of }\mathbb{R}^{d}\},~{}d\in\mathbb{N},$
such that the following requirements are satisfied:
  1. (1)
  2. (2)
    For each $n\in\mathbb{N}$, $\mathscr{S}_{n}$ contains the diagonals $\{(x_{1},\dots,x_{n}):x_{i}=x_{j}\},$ for all $1\leq i<j\leq n$.
  3. (3)
  4. (4)
    For each $n\in\mathbb{N}$, if $A\in\mathscr{S}_{n+1}$, then $\pi(A)\in\mathscr{S}_{n}$, where $\pi:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ is the projection to the first $n$ coordinates.
  5. (5)
Given a structure $\mathscr{S}$, we say that $A\subseteq\mathbb{R}^{d}$ is definable if $A$ belongs to $\mathscr{S}_{d}$. We say that a function $f:B\to\mathbb{R}^{m}$ is definable, for $B\subseteq\mathbb{R}^{d}$, if its graph $\{(x,f(x):x\in B\}$ belongs to $\mathscr{S}_{d+m}$.
We note the following basic properties, which are proved directly from Definition B.1
Lemma B.2.
Let $\mathscr{S}$ be a structure on the field of real numbers. Then:
  1. (1)
    The restriction of a definable function to a definable set is definable function on that set.
  2. (2)
    If $f:\mathbb{R}^{d}\to\mathbb{R}$ is definable, and $A\subseteq\mathbb{R}^{d}$, is definable, then the subsets:
    $\left\{\mathbf{x}\in A\mid f(\mathbf{x})=0\right\},\ \left\{\mathbf{x}\in A\mid f(\mathbf{x})>0\right\}$
    are definable.
  3. (3)
    If $f:A\to B$ and $g:B\to C$ are definable, then the composition $g\circ f$ is also definable.
  4. (4)
    If $f,g:A\to\mathbb{R}$ are definable, then $f+g,$ $f-g,$ $f\cdot g$ are definable, and $f/g$ on the domain $\{x\in A:g(x)\neq 0\}$ is definable.
The notion of a structures belongs to model theory, and it is most natural to describe definable sets using first order logic formulae, see [Geom_cat_and_o_min_str, Appendix A] for a concise introduction. The advantage of the definition above is in allowing to define structures without resorting to notions of model theory.
The proof of the statement below demonstrates the technique of proving the definability of functions using first-order logic formulae. The claim is useful for us since the type of the function considered there is used in the proof of Theorem 3.2.
Lemma B.3.
Suppose that $f:\mathbb{R}^{2}\to\mathbb{R}$ is a bounded definable function, and let $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ be definable functions such that $a(s)<b(s),~{}\forall s$. Then the function $\Psi:\mathbb{R}\to\mathbb{R}$ defined by
$\Psi(s):=\sup\left\{|f(s,t)|:t\in\left[a(s),b(s)\right]\right\},\ \forall s\in\mathbb{R},$
is definable.
Proof.
To prove the $\Psi$ is definable, we need to verify that the graph of $\Psi$, $\Gamma(\Psi):=\{(s,\Psi(s)):s\in\mathbb{R}\}$ is definable. Consider the definable set:
$A_{1}:=\{(s,y,t):a(s)\leq t\leq b(s)\text{ and }|f(s,t)|>y\},$
and consider the projection $\pi_{3,2}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ to the first coordinates. Then $B_{1}:=(\pi_{2,3}(A_{1}))^{c}$ is definable, and it can be alternatively described by:
$B_{1}:=\{(s,y):(\forall t\in[a(s),b(s)])\ |f(s,t)|\leq y\}.$
In a similar way (while slightly complicated), the set:
$B_{2}:=\{(s,y):(\forall\epsilon>0)(\exists t\in[a(s),b(s)])\ y-\epsilon\leq|f(s,t)|\}$
is definable. The graph of $\Psi$ is then $B_{1}\cap B_{2}$. ∎
B.1. O-minimal structures
The notion of o-minimal structures was introduced in [generalization_tarski_seind_dries], as a generalization of the real semi-algebraic geometry.
Definition B.4.
(o-minimal structures) A structure $\mathscr{S}$ expanding the real field is called o-minimal if $\mathscr{S}_{1}$ consists of all possible sets which are finite unions of points and intervals.
The semi-algebraic sets are definable in every o-minimal structure. Here, the semi-algebraic sets are boolean combinations of sets of the form $\{x\in\mathbb{R}^{d}:P(x)=0\}$, $\{x\in\mathbb{R}^{d}:P(x)>0\}$, where $P\in\mathbb{R}[X_{1},\dots,X_{d}]$. By the Tarski-Seidenberg theorem, the semi-algebraic sets form an o-minimal structure.
We collect the following important results, see [van_den_dries_omin_book] for detailed proofs. For the following, fix an o-minimal structure.
Theorem B.5.
[van_den_dries_omin_book, Ch. 3, Thm. 1.2 and Ch. 7, Prop. 2.5] Suppose that $f:(a,b)\to\mathbb{R}$ is definable, where $-\infty\leq a<b\leq\infty$. Then, there exist $a_{0},a_{1},...,a_{k+1}$ such that $a=:a_{0}<a_{1}<\dots<a_{k}<a_{k+1}:=b$ and such that in each sub-interval $I=(a_{i},a_{i+1})$, $f|_{I}$ is differentiable and $f|_{I}$ is either strictly monotonic, or constant.
A direct consequence of the monotonicity theorem is that a definable function $f:\mathbb{R}\to\mathbb{R}$ either converges as $t\to\infty$ or diverges to $\infty$ or to $\infty$- ∞.
Theorem B.6.
[van_den_dries_omin_book, Ch.3, Cor. 3.7]. If $A\subseteq\mathbb{R}^{m}$ is definable, than $A$ has finitely many connected components. Moreover, the following uniformity holds. For $A\subseteq\mathbb{R}^{d}\times\mathbb{R}^{n}$ and $x_{0}\in\mathbb{R}^{d}$, consider
$A_{x_{0}}:=\{(x_{0},y):(x_{0},y)\in A\}.$
Then, for a definable $A$, there exists an $N\in\mathbb{N}$ such for all $x\in\mathbb{R}^{m}$, the set $A_{x}$ has at most $N$ connected components.
Theorem B.7.
[van_den_dries_omin_book, Ch. 3, Prop. 1.2]. Suppose that $A\subseteq\mathbb{R}^{m+n}$ is definable. Let $\pi_{m+n,m}:\mathbb{R}^{m+n}\to\mathbb{R}^{m}$ be the projection to the first $m$ coordintates. Then, there is a definable function $f:\pi_{m+n,m}(A)\to\mathbb{R}^{n}$ such that the graph of $f$ is contained in $A$.